OpenStudy (anonymous):

Multiply (4 + i)(1 – 5i). i got the answer I just want to see if I am correct.

OpenStudy (anonymous):

@Michele_Laino

Nnesha (nnesha):

please post your answer with the `WORK` :=)) i'll check it

OpenStudy (anonymous):

(4+i)*(1-5i) 4-20i+1i-5i^2 4-19i-5i^2 4-19i-5i^2 is my Answer

OpenStudy (breyliance):

im sure your right i think

Nnesha (nnesha):

alright very nice!!! \[i=\sqrt{-1}\] so i^2 = -1 so replace i^2 with -1 :=))

OpenStudy (anonymous):

ok thanks, My answers are 9-19i -1-19i 4-5i^2 4-19i-5i^2 so the last one is correct right

Nnesha (nnesha):

hmm no use the fact `i = sqrt{-1}` so i^2 = -1 substitute i^2 for -1 :=))

Nnesha (nnesha):

\[i=\sqrt{-1} ~~~~~ i^2=(\sqrt{-1})^2 \rightarrow i^2=-1\] i is an imaginary number that is equal to sqrt{-1} when you take square of it you willget i^2 = -1

OpenStudy (anonymous):

ok can you walk me through the steps then? because now Im a bit confused xD

OpenStudy (anonymous):

the last steps after what I did? Im not very good with this :/

Nnesha (nnesha):

okay your steps are correct i'll take it from 4-19i-5i^2 \[\huge\rm 4-19i -5\color{ReD}{i^2}\] i'm going to replace i^2 with -1 using the fact \[\large\rm i =\sqrt{-1} ~~and~~ i^2=-1\] \[\huge\rm 4-19i -5\color{ReD}{(-1)}\] now simplify

OpenStudy (anonymous):

4-19i-5(-1) -1-19i(-1) -2-19i <- that is it?

Nnesha (nnesha):

mm no first multiply -5 times -1 and then combine like terms

Nnesha (nnesha):